# Multiple Proposer Transaction Fee Mechanism Design: Robust Incentives Against Censorship and Bribery

### v1

@@ 5 Single TFM
The system splits the fee as follows: after subtracting the burning fee, the block producer receives an amount that is equal to the cost they incur for processing each transaction - namely $\mu^{Cost}_{BP}\cdot s$. If the remaining amount after burning is less than $\mu^{Cost}_{BP}\cdot s$, the block producer receives the entire remaining amount. Any residual amount beyond this is then distributed as follows: $z$ fraction of this amount is allocated to the committee and $(1-z)$ fraction to the block producer. In Appendix C.5, we examine how $z$ affects the minimum bribe needed for a briber to make a transaction be omitted from a block.
@@ Appendix B Proofs for Double TFM
    $\min\{v_{t},r+\mu^{Cost}_{BP}\}=r+\mu^{Cost}_{BP}$ : The current utility of a user who bids $b_{t}=(\delta^{CM}_{t},\delta^{BP}_{t},c_{t})=(0,\mu^{Cost}_{BP},\min\{v_{t},r+\mu^{Cost}_{BP}\})$ in this case is $v_{t}-r-\mu^{Cost}_{BP}\geq 0$ regardless of their beliefs, because their transaction will be included in the block by the block producer irrespective of the bids of the other users and the type of the block producer. The utility of the user when their transaction is not included in the block is 00. Thus, the user can increase their utility only if they can set a lower $c_{t}$ and make their transaction still included in the block. The burning fee is $r$ which means that regardless of their choice of $\delta^{CM}_{t},\delta^{BP}_{t}$, if they decrease $c_{t}$, the block producer fee becomes lower than $\mu^{Cost}_{BP}\cdot s$. This means that this transaction will not be included in the block.
@@ B.2 Myopic Committee Bayesian-Nash Incentive Compatible (MCBN)
    • When this fake transaction has a committee fee that makes another includer with a smaller order prefer it over another transaction $t^{\prime}$, then their utility can be affected if they “steal” $t^{\prime}$. However, their utility cannot increase because (i) the maximum includer $j$ can gain by this deviation is the committee fee of the omitted transaction $t^{\prime}$ and (ii) includer $j$ needs to pay the committee fee, the block producer fee and the burning fee of this omitted transaction $t^{\prime}$ This means that the utility of the includer will decrease.
@@ B.3 Myopic Block Producer Bayesian-Nash Incentive Compatible (MBBN)
    • If $size_{L_{BP}}\geq c_{block}+1$: The most profitable deviation the block producer can perform is to replace transaction $t_{0}$ with the transaction that has position $c_{block}+1$ in $L_{BP}$ If they do so, they will gain the bribe loss $f_{BP}-\max\{f_{c_{block}+1,BP},\mu^{Cost}_{BP}\cdot s\}=f_{BP}-f_{c_{block}+1,BP}$ but they will lose $f_{BP}-f_{c_{block}+1,BP}$ because $t_{0}$ has a higher block producer fee. Moreover, the block producer incurs the same cost $\mu^{Cost}_{BP}\cdot s$ for both $t_{0}$ and the transaction with position $c_{block}+1$ in $L_{BP}$ This means that their utility will not increase.
@@
    • They can omit transaction $t_{0}$ (losing its fee $f_{BP}$) and add no fake transaction. This means that they will lose their entire block rewards. Their utility in this case becomes 00, which is not higher than their current utility if the bribe loss is no higher than $sum_{max,c_{block}}$.
@@ C.2 Myopic Committee Bayesian-Nash Incentive Compatible (MCBN)
  • In order for the fake transaction to affect which transactions the other includers or the block producer include in their inclusion list and block respectively, they need to have a fee that will be paid by includer $j$.”
@@
    • When this fake transaction gives a committee fee that makes another includer with a smaller order prefer it over another transaction $t^{\prime}$, then their utility can be affected if they “steal” $t^{\prime}$. However, their utility cannot increase because (i) the maximum includer $j$ can gain by this deviation is the committee fee of the omitted transaction $t^{\prime}$ and (ii) includer $j$ needs to pay the committee fee of this omitted transaction $t^{\prime}$, the burning fee, $\mu^{Cost}_{BP}$, and $(1-z)/z$ times the committee fee of $t^{\prime}$ (this is the fee that is awarded to the block producer apart from $\mu^{Cost}_{BP}$). This means that the utility of includer $j$ will decrease.
@@ C.5 Intuition about How the Choice $z,c_{t_{0}}$ Affect the Minimum Bribe Needed for Censorship
  • When $[\frac{r}{m}>c_{t_{0}}-r]$ OR $[\frac{r}{m}\leq c_{t_{0}}-r\text{ and }0\leq z\leq\frac{\frac{r}{m}}{(c_{t_{0}}-r)}]$ the equation $2,3$ is equal to
@@ Appendix E Theorems and Proofs for Simplified Model
  • Let $sum$ be the block rewards (block producer fees minus costs).

### v2

@@ 5 Single TFM
The system splits the fee as follows: after subtracting the burning fee, the block producer receives an amount that is equal to the cost they incur for processing each transaction - namely $\mu^{Cost}_{BP}\cdot s$. If the remaining amount after burning is less than $\mu^{Cost}_{BP}\cdot s$, the block producer receives the entire remaining amount. Any residual amount beyond this is then distributed as follows: $z$ fraction of this amount is allocated to the committee and (1z)(1-z) fraction to the block producer. In Appendix C.5, we examine how $z$ affects the minimum bribe needed for a briber to make a transaction be omitted from a block.
@@ Appendix B Proofs for Double TFM
    $\min\{v_{t},r+\mu^{Cost}_{BP}\}=r+\mu^{Cost}_{BP}$ : The current utility of a user who bids $b_{t}=(\delta^{CM}_{t},\delta^{BP}_{t},c_{t})=(0,\mu^{Cost}_{BP},\min\{v_{t},r+\mu^{Cost}_{BP}\})$ in this case is $v_{t}-r-\mu^{Cost}_{BP}\geq 0$ regardless of their beliefs, because their transaction will be included in the block by the block producer irrespective of the bids of the other users and the type of the block producer. The utility of the user when their transaction is not included in the block is 0. Thus, the user can increase their utility only if they can set a lower $c_{t}$ and make their transaction still included in the block. The burning fee is $r$ which means that regardless of their choice of $\delta^{CM}_{t},\delta^{BP}_{t}$, if they decrease $c_{t}$, the block producer fee becomes lower than $\mu^{Cost}_{BP}\cdot s$. This means that this transaction will not be included in the block.
@@ B.2 Myopic Committee Bayesian-Nash Incentive Compatible (MCBN)
    • When this fake transaction has a committee fee that makes another includer with a smaller order prefer it over another transaction $t^{\prime}$, then their utility can be affected if they “steal” $t^{\prime}$. However, their utility cannot increase because (i) the maximum includer $j$ can gain by this deviation is the committee fee of the omitted transaction $t^{\prime}$ and (ii) includer $j$ needs to pay the committee fee, the block producer fee and the burning fee of this omitted transaction $t^{\prime}$ . This means that the utility of the includer will decrease.
@@ B.3 Myopic Block Producer Bayesian-Nash Incentive Compatible (MBBN)
    • If $size_{L_{BP}}\geq c_{block}+1$: The most profitable deviation the block producer can perform is to replace transaction $t_{0}$ with the transaction that has position $c_{block}+1$ in $L_{BP}$ . If they do so, they will gain the bribe loss $f_{BP}-\max\{f_{c_{block}+1,BP},\mu^{Cost}_{BP}\cdot s\}=f_{BP}-f_{c_{block}+1,BP}$ but they will lose $f_{BP}-f_{c_{block}+1,BP}$ because $t_{0}$ has a higher block producer fee. Moreover, the block producer incurs the same cost $\mu^{Cost}_{BP}\cdot s$ for both $t_{0}$ and the transaction with position $c_{block}+1$ in $L_{BP}$ . This means that their utility will not increase.
@@
    • They can omit transaction $t_{0}$ (losing its fee $f_{BP}$) and add no fake transaction. This means that they will lose their entire block rewards. Their utility in this case becomes 0, which is not higher than their current utility if the bribe loss is no higher than $sum_{max,c_{block}}$.
@@ C.2 Myopic Committee Bayesian-Nash Incentive Compatible (MCBN)
  • In order for the fake transaction to affect which transactions the other includers or the block producer include in their inclusion list and block respectively, they need to have a fee that will be paid by includer jj.”
@@
    • When this fake transaction gives a committee fee that makes another includer with a smaller order prefer it over another transaction $t^{\prime}$, then their utility can be affected if they “steal” $t^{\prime}$. However, their utility cannot increase because (i) the maximum includer $j$ can gain by this deviation is the committee fee of the omitted transaction $t^{\prime}$ and (ii) includer $j$ needs to pay the committee fee of this omitted transaction $t^{\prime}$, the burning fee, $\mu^{Cost}_{BP}$, and (1z)/z(1-z)/z times the committee fee of $t^{\prime}$ (this is the fee that is awarded to the block producer apart from $\mu^{Cost}_{BP}$). This means that the utility of includer $j$ will decrease.
@@ C.5 Intuition about How the Choice $z,c_{t_{0}}$ Affect the Minimum Bribe Needed for Censorship
[…]
@@ Appendix E Theorems and Proofs for Simplified Model
  • Let sumsum be the block rewards (block producer fees minus costs).